# Accelerated Mirror Descent for Non-Euclidean Star-convex Functions

### v1

@@ Appendix B Useful lemma for convergence proof
Lemma B.1.
(Young’s inequality [33]) Let $a,b\geq 0$ and $p,q\geq 1$ such that $\frac{1}{p}+\frac{1}{q}=1$. Then it holds, that
@@
Let $f$ be a convex function and $\nu$ be convex and continuously differentiable. For
@@
As $u^{*}$ is a minimizer, it fulfills for all $u\in{\cal X}$ the first-order optimality condition
@@
with the gradient taken with respect to the first entry. Applying the three-points identity [8], yields
@@
Using Young’s inequality B.1 with $\frac{1}{a}+\frac{1}{b}=1$, leads to $t\leq\frac{1}{az}t^{a}+\frac{1}{b}z^{b-1}.$ By choosing $t:=\frac{\|x-y\|^{\kappa}}{\kappa}$, $a=\frac{q}{\kappa}$, $b=\frac{q}{q-\kappa}$, $z=(b\delta)^{\frac{1}{b-1}}=(b\delta)^{r}$, it follows
@@
Our proof can be decomposed in three major steps. We first derive a bound for $F(x_{t}^{md})-F_{\star}$ by analysing our mirror descent step, then we proceed an acceleration step comparing $F(x_{t+1}^{ag})$ and $F(x_{t}^{md})$ and last we derive the parameters satisfying the different conditions needed.
@@ C.1 Mirror descent analysis
$\eta_{t}\langle\nabla F(x_{t}^{md}),x_{t+1}\rangle+D^{\psi}(x_{t+1},x_{t})\leq\eta_{t}\langle\nabla F(x_{t}^{md}),x\rangle+D^{\psi}(x,x_{t})-D^{\psi}(x,x_{t+1}).$
@@
$\eta_{t}\langle\nabla F(x_{t}^{md}),x_{t+1}-x_{\star}\rangle+D^{\psi}(x_{t+1},x_{t})\leq D^{\psi}(x_{\star},x_{t})-D^{\psi}(x_{\star},x_{t+1}).$
Decomposing $\eta_{t}\langle\nabla F(x_{t}^{md}),x_{t+1}-x_{\star}\rangle$ in
@@
along with
$\frac{\eta_{t}}{\tau}\Big{(}F(x_{t}^{md})-F_{\star}\Big{)}\leq\eta_{t}\langle\nabla F(x_{t}^{md}),x_{t}^{md}-x_{\star}\rangle,$
following from the the star-convexity of $F$, gives
$\frac{\eta_{t}}{\tau}\Big{(}F(x_{t}^{md})-F_{\star}\Big{)}+\eta_{t}\langle\nabla F(x_{t}^{md}),x_{t+1}-x_{t}^{md}\rangle+D^{\psi}(x_{t+1},x_{t})\leq D^{\psi}(x_{\star},x_{t})-D^{\psi}(x_{\star},x_{t+1}).$
@@ C.2 Choice of aggregated point
${}\leq{}\frac{\mu C^{q_{\ast}}\eta_{t}^{q_{\ast}}}{q_{\ast}\alpha_{t}^{q_{\ast}}}\|x_{t+1}^{ag}-x_{t}^{md}\|^{(q-1)q_{\ast}}$
${}={}\frac{\mu C^{q_{\ast}}\eta_{t}^{q_{\ast}}}{q_{\ast}\alpha_{t}^{q_{\ast}}}\|x_{t+1}^{ag}-x_{t}^{md}\|^{q},$
[…]
By setting $u=x_{t}^{md}$, we obtain from first-order optimality, that
@@
$\leq$ $\frac{\mu C^{q_{\ast}}\eta_{t}^{q_{\ast}}}{\alpha_{t}^{q_{\ast}}}\frac{\|x_{t+1}^{ag}-x_{t}^{md}\|^{q}}{q_{\ast}}$
$\leq$ $\frac{C^{q_{\ast}}\eta_{t}^{q_{\ast}}}{\alpha_{t}^{q_{\ast}}}\Bigg{(}\alpha_{t}\Big{(}F(x_{t}^{md})-F(x_{t+1}^{ag})\Big{)}+\frac{M^{1/r}}{\mu^{1/r}}\Big{(}L\alpha_{t}\Big{)}^{(r+1)/r}\Bigg{)}$
$=$ $\frac{C^{q_{\ast}}\eta_{t}^{q_{\ast}}}{\alpha_{t}^{q_{\ast}-1}}\Big{(}F(x_{t}^{md})-F(x_{t+1}^{ag})\Big{)}+\frac{C^{q_{\ast}}\eta_{t}^{q_{\ast}}}{\alpha_{t}^{q_{\ast}}}\frac{M^{1/r}}{\mu^{1/r}}\Big{(}L\alpha_{t}\Big{)}^{(r+1)/r}$
@@ C.3 Simplified equation
$A_{t}$ $:=\frac{C^{q_{\ast}}\eta_{t}^{q_{\ast}}}{\alpha_{t}^{q_{\ast}-1}}$ (15)
$B_{t}$ $:=\frac{C^{q_{\ast}}\eta_{t}^{q_{\ast}}}{\alpha_{t}^{q_{\ast}}}\frac{M^{1/r}}{\mu^{1/r}}\Big{(}L\alpha_{t}\Big{)}^{1+1/r}=\frac{C^{q_{\ast}}\eta_{t}^{q_{\ast}}}{\alpha_{t}^{q_{\ast}-1-1/r}}\frac{M^{1/r}}{\mu^{1/r}}L^{1+1/r}.$
Rearranging (11) and combining it with equation (14) from the previous subsections leads to
@@
$\leq$ $D^{\psi}(x_{\star},x_{t})-D^{\psi}(x_{\star},x_{t+1})+\eta_{t}\langle\nabla F(x_{t}^{md}),x_{t}^{md}-x_{t}\rangle$
@@
With the binary search on $\lambda_{t}$, Algorithm 1 achieves
@@
$A_{t}\Big{(}F(x_{t+1}^{ag})-F_{\star}\Big{)}\leq$ $D^{\psi}(x_{\star},x_{t})-D^{\psi}(x_{\star},x_{t+1})+\eta_{t}\epsilon_{t}$ (17)
@@
$\begin{cases}A_{t}-\frac{\eta_{t}}{\tau}&\leq A_{t-1}\\ B_{t}&\leq\alpha^{\frac{q}{q-\kappa}}\frac{C^{q_{\ast}}}{t}\frac{M^{1/r}}{\mu^{1/r}}L^{1+1/r}.\end{cases}$
@@
$A_{T}\Big{(}F(x_{T+1}^{ag})-F_{\star}\Big{)}\leq D^{\psi}(x_{\star},x_{1})+\mathcal{O}\Big{(}\alpha^{\frac{q}{q-\kappa}}\frac{C^{q_{\ast}}L^{1+1/r}}{\mu^{1/r}}\log(T)\Big{)}.$ (18)
@@
$\beta=q-\kappa+\frac{q-\kappa}{q},\quad\gamma\geq 0,\quad\Big{(}\frac{\eta_{t}}{\alpha_{t}}\Big{)}^{q_{\star}-1}=\frac{q-\beta}{\tau C^{q_{\ast}}}t^{\gamma(q_{\star}-1)}$
and recall that $1+\frac{1}{r}=1+\frac{\kappa}{q-\kappa}=\frac{q}{q-\kappa}$ and $\frac{q_{\ast}}{q}=\frac{1}{q-1}$. Then it follows
$B_{t}\leq\alpha^{\frac{q}{q-\kappa}}\frac{C^{q_{\ast}}}{t}\frac{M^{1/r}}{\mu^{1/r}}L^{1+1/r},$
@@
$\frac{C^{q_{\ast}}\eta_{t}^{q_{\ast}}}{\alpha_{t}^{q_{\ast}-1-1/r}}\frac{M^{1/r}}{\mu^{1/r}}L^{1+1/r}\leq\alpha^{\frac{q}{q-\kappa}}\frac{C^{q_{\ast}}}{t}\frac{M^{1/r}}{\mu^{1/r}}L^{1+1/r}$
@@
$\alpha_{t}^{1+1/r}\leq\frac{\alpha^{\frac{q}{q-\kappa}}}{t^{\gamma q_{\ast}+1}}\Leftrightarrow$ $\quad\alpha_{t}\leq\frac{\alpha}{t^{\frac{\gamma(q-\kappa)}{q-1}+\frac{q-\kappa}{q}}}.$
@@
$A_{t}-A_{t-1}=C^{q_{\ast}}\Big{(}\frac{\eta_{t}^{q_{\ast}}}{\alpha_{t}^{q_{\ast}-1}}-\frac{\eta_{t-1}^{q_{\ast}}}{\alpha_{t-1}^{q_{\ast}-1}}\Big{)}=(q-\beta)C^{q_{\ast}}\left(\eta_{t}\Big{(}\frac{\eta_{t}}{\alpha_{t}}\Big{)}^{q_{\star}-1}-\eta_{t-1}\Big{(}\frac{\eta_{t-1}}{\alpha_{t-1}}\Big{)}^{q_{\star}-1}\right)\leq\frac{\eta_{t}}{\tau}$
and therefore
$(q-\beta)\Big{(}t^{\gamma(q_{\ast}-1)}-\frac{\eta_{t-1}}{\eta_{t}}(t-1)^{\gamma(q_{\ast}-1)}\Big{)}\leq 1.$
@@
$\alpha_{t}$ $:=\frac{\alpha}{t^{\beta}}$
$\eta_{t}$ $:=\alpha_{t}\Big{(}\frac{q-\beta}{\tau C^{q_{\ast}}}t^{\gamma(q_{\star}-1)}\Big{)}^{\frac{1}{q_{\star}-1}}=\alpha_{t}\Big{(}\frac{q-\beta}{\tau C^{q_{\ast}}}t\Big{)}^{q-1}.$
@@
$\frac{A_{t}-A_{t-1}}{\eta_{t}}=$ $\frac{q-\beta}{\tau}\Big{(}t-\frac{\eta_{t-1}}{\eta_{t}}(t-1)\Big{)}$
$=$ $\frac{q-\beta}{\tau}\frac{1}{t^{q-1-\beta}}\Big{(}t^{q-\beta}-(t-1)^{q-\beta}\Big{)}$
@@
By choosing $\epsilon_{t}:=\frac{B_{t}}{\eta_{t}}$ such that $\eta_{t}\epsilon_{t}\leq B_{t}$, leads to
$A_{t}$ $=\frac{C^{q_{\ast}}\eta_{t}^{q_{\ast}}}{\alpha_{t}^{q_{\ast}-1}}=C^{q_{\ast}}\Big{(}\frac{\eta_{t}}{\alpha_{t}}\Big{)}^{q_{\ast}}\alpha_{t}=C^{q_{\ast}}\Big{(}\frac{q-\beta}{\tau C^{q_{\ast}}}t\Big{)}^{\frac{q_{\star}}{q_{\star}-1}}\alpha_{t}$
$\propto C^{q_{\ast}}\cdot\Big{(}\frac{t}{C^{q_{\ast}}\tau}\Big{)}^{q}\cdot\frac{\alpha}{t^{\beta}}=C^{q_{\ast}(1-q)}\cdot\alpha t^{q-\beta}\cdot\frac{1}{\tau^{q}}$
$A_{t}$ $\propto\frac{\alpha}{\tau^{q}C^{q}}t^{\frac{\kappa q+\kappa-q}{q}}$
@@
$F(x_{T+1}^{ag})-F_{\star}$ $=\mathcal{O}_{q,\kappa}\Big{(}\frac{C^{q}\tau^{q}D^{\psi}(x_{\star},x_{1})}{\alpha T^{\frac{\kappa q+\kappa-q}{q}}}+\frac{\alpha^{\frac{\kappa}{q-\kappa}}C^{q_{\ast}}L^{\frac{q}{q-\kappa}}}{\mu^{\frac{\kappa}{q-\kappa}}}\frac{C^{q}\tau^{q}\log(T)}{T^{\frac{\kappa q+\kappa-q}{q}}}\Big{)},$
@@
Appendix D Proof for binary search
Lemma D.1.

### v2

@@ Appendix B Useful lemma for convergence proof
Lemma B.1 (Young’s inequality, [young]).
Let $a,b\geq 0$ and $p,q\geq 1$ such that $\frac{1}{p}+\frac{1}{q}=1$. Then it holds, that
@@
Let $f$, $\nu$ be convex functions and let $\nu$ be continuously differentiable. For
@@
As $u^{\star}$ is a minimizer, it fulfills for all $u\in{\cal X}$ the first-order optimality condition
@@
with the gradient of the divergence $D^{\nu}$ taken with respect to the first argument. Applying the three-points identity [CT93], yields
@@
Using Young’s inequality B.1 with $\frac{1}{a}+\frac{1}{b}=1$, leads to $t\leq\frac{1}{az}t^{a}+\frac{1}{b}z^{b-1}.$ By choosing $t:=\frac{\|x-y\|^{\kappa}}{\kappa}$, $a=\frac{q}{\kappa}$, $b=\frac{q}{q-\kappa}$, $z=(b\delta)^{\frac{1}{b-1}}=(b\delta)^{r}$, it follows that
@@
Our proof can be decomposed in three major steps. We first derive a bound for $F(x_{t}^{md})-F_{\star}$ by analyzing our mirror descent step, then we proceed with an acceleration step comparing $F(x_{t+1}^{ag})$ and $F(x_{t}^{md})$ and finally we tune the parameters to satisfy the required conditions.
@@ C.1 Mirror descent analysis
$\eta_{t}\langle\nabla F(x_{t}^{md}),x_{t+1}\rangle+D_{\psi}(x_{t+1},x_{t})\leq\eta_{t}\langle\nabla F(x_{t}^{md}),x\rangle+D_{\psi}(x,x_{t})-D_{\psi}(x,x_{t+1}).$
@@
$\eta_{t}\langle\nabla F(x_{t}^{md}),x_{t+1}-x_{\star}\rangle+D_{\psi}(x_{t+1},x_{t})\leq D_{\psi}(x_{\star},x_{t})-D_{\psi}(x_{\star},x_{t+1}).$
We split
@@
and use star-convexity of $F$
$\frac{\eta_{t}}{\tau}\Big{(}F(x_{t}^{md})-F_{\star}\Big{)}\leq\eta_{t}\langle\nabla F(x_{t}^{md}),x_{t}^{md}-x_{\star}\rangle$
to find
$\frac{\eta_{t}}{\tau}\Big{(}F(x_{t}^{md})-F_{\star}\Big{)}+\eta_{t}\langle\nabla F(x_{t}^{md}),x_{t+1}-x_{t}^{md}\rangle+D_{\psi}(x_{t+1},x_{t})\leq D_{\psi}(x_{\star},x_{t})-D_{\psi}(x_{\star},x_{t+1}).$
@@ C.2 Choice of aggregated point
${}={}\frac{\mu\eta_{t}^{q_{\ast}}}{q_{\ast}\alpha_{t}^{q_{\ast}}}\|x_{t+1}^{ag}-x_{t}^{md}\|^{(q-1)q_{\ast}}$
${}={}\frac{\mu\eta_{t}^{q_{\ast}}}{q_{\ast}\alpha_{t}^{q_{\ast}}}\|x_{t+1}^{ag}-x_{t}^{md}\|^{q},$
where the penultimate equality follows from Lemma A.2. By setting $u=x_{t}^{md}$, we obtain from first-order optimality, that
@@
$\leq$ $\frac{\mu\eta_{t}^{q_{\ast}}}{\alpha_{t}^{q_{\ast}}}\frac{\|x_{t+1}^{ag}-x_{t}^{md}\|^{q}}{q_{\ast}}$
$\leq$ $\frac{\eta_{t}^{q_{\ast}}}{\alpha_{t}^{q_{\ast}}}\Bigg{(}\alpha_{t}\Big{(}F(x_{t}^{md})-F(x_{t+1}^{ag})\Big{)}+\frac{M^{1/r}}{\mu^{1/r}}\Big{(}L\alpha_{t}\Big{)}^{(r+1)/r}\Bigg{)}$
$=$ $\frac{\eta_{t}^{q_{\ast}}}{\alpha_{t}^{q_{\ast}-1}}\Big{(}F(x_{t}^{md})-F(x_{t+1}^{ag})\Big{)}+\frac{\eta_{t}^{q_{\ast}}}{\alpha_{t}^{q_{\ast}}}\frac{M^{1/r}}{\mu^{1/r}}\Big{(}L\alpha_{t}\Big{)}^{(r+1)/r}$
@@ C.3 Simplified equation
$A_{t}$ $:=\frac{\eta_{t}^{q_{\ast}}}{\alpha_{t}^{q_{\ast}-1}}$ (15)
$B_{t}$ $:=\frac{\eta_{t}^{q_{\ast}}}{\alpha_{t}^{q_{\ast}}}\frac{M^{1/r}}{\mu^{1/r}}\Big{(}L\alpha_{t}\Big{)}^{1+1/r}=\frac{\eta_{t}^{q_{\ast}}}{\alpha_{t}^{q_{\ast}-1-1/r}}\frac{M^{1/r}}{\mu^{1/r}}L^{1+1/r}.$
Rearranging (11) and combining it with equation (14) leads to
@@
$\leq$ $D_{\psi}(x_{\star},x_{t})-D_{\psi}(x_{\star},x_{t+1})+\eta_{t}\langle\nabla F(x_{t}^{md}),x_{t}^{md}-x_{t}\rangle$
@@
With the binary search for $\lambda_{t}$, Algorithm 1 achieves
@@
$A_{t}\Big{(}F(x_{t+1}^{ag})-F_{\star}\Big{)}\leq$ $D_{\psi}(x_{\star},x_{t})-D_{\psi}(x_{\star},x_{t+1})+\eta_{t}\epsilon_{t}$ (17)
@@
$\begin{cases}A_{t}-\frac{\eta_{t}}{\tau}&\leq A_{t-1}\\ B_{t}&\leq\alpha^{\frac{q}{q-\kappa}}\frac{1}{t}\frac{M^{1/r}}{\mu^{1/r}}L^{1+1/r}.\end{cases}$
@@
$A_{T}\Big{(}F(x_{T+1}^{ag})-F_{\star}\Big{)}\leq D_{\psi}(x_{\star},x_{1})+\mathcal{O}\Big{(}\alpha^{\frac{q}{q-\kappa}}\frac{L^{1+1/r}}{\mu^{1/r}}\log(T)\Big{)}.$ (18)
@@
$\beta=q-\kappa+\frac{q-\kappa}{q},\quad\gamma\geq 0,\quad\Big{(}\frac{\eta_{t}}{\alpha_{t}}\Big{)}^{q_{\star}-1}=\frac{1}{\tau(q-\beta)}t^{\gamma(q_{\star}-1)}$
and recall that $q-\beta=\frac{q\kappa+\kappa-q}{q}>0$, $1+\frac{1}{r}=1+\frac{\kappa}{q-\kappa}=\frac{q}{q-\kappa}$ and $\frac{q_{\ast}}{q}=\frac{1}{q-1}$. Then it follows
$B_{t}\leq\alpha^{\frac{q}{q-\kappa}}\frac{1}{t}\frac{M^{1/r}}{\mu^{1/r}}L^{1+1/r},$
@@
$\frac{\eta_{t}^{q_{\ast}}}{\alpha_{t}^{q_{\ast}-1-1/r}}\frac{M^{1/r}}{\mu^{1/r}}L^{1+1/r}\leq\alpha^{\frac{q}{q-\kappa}}\frac{1}{t}\frac{M^{1/r}}{\mu^{1/r}}L^{1+1/r}$
@@
$\alpha_{t}^{1+1/r}\leq\Big{(}\tau(q-\beta)\Big{)}^{\frac{q_{\star}}{q_{\star}-1}}\frac{\alpha^{\frac{q}{q-\kappa}}}{t^{\gamma q_{\ast}+1}}\Leftrightarrow$ $\quad\alpha_{t}\leq\Big{(}\tau(q-\beta)\Big{)}^{q-\kappa}\frac{\alpha}{t^{\frac{\gamma(q-\kappa)}{q-1}+\frac{q-\kappa}{q}}}.$
@@
$A_{t}-A_{t-1}=\Big{(}\frac{\eta_{t}^{q_{\ast}}}{\alpha_{t}^{q_{\ast}-1}}-\frac{\eta_{t-1}^{q_{\ast}}}{\alpha_{t-1}^{q_{\ast}-1}}\Big{)}=\left(\eta_{t}\Big{(}\frac{\eta_{t}}{\alpha_{t}}\Big{)}^{q_{\star}-1}-\eta_{t-1}\Big{(}\frac{\eta_{t-1}}{\alpha_{t-1}}\Big{)}^{q_{\star}-1}\right)\leq\frac{\eta_{t}}{\tau}$
is equialent to
$\frac{A_{t}-A_{t-1}}{\eta_{t}}=\frac{1}{\tau(q-\beta)}\Big{(}t^{\gamma(q_{\ast}-1)}-\frac{\eta_{t-1}}{\eta_{t}}(t-1)^{\gamma(q_{\ast}-1)}\Big{)}\leq\frac{1}{\tau}.$
@@
$\alpha_{t}$ $:=\Big{(}\tau(q-\beta)\Big{)}^{q-\kappa}\frac{\alpha}{t^{\beta}}$
$\eta_{t}$ $:=\alpha_{t}\Big{(}\frac{1}{\tau(q-\beta)}t^{\gamma(q_{\star}-1)}\Big{)}^{\frac{1}{q_{\star}-1}}=\alpha_{t}\Big{(}\frac{t}{\tau(q-\beta)}\Big{)}^{q-1}.$
@@
$\frac{A_{t}-A_{t-1}}{\eta_{t}}=$ $\frac{1}{\tau(q-\beta)}\Big{(}t-\frac{\eta_{t-1}}{\eta_{t}}(t-1)\Big{)}$
$=$ $\frac{1}{\tau(q-\beta)}\frac{1}{t^{q-1-\beta}}\Big{(}t^{q-\beta}-(t-1)^{q-\beta}\Big{)}$
@@
By choosing $\epsilon_{t}:=\frac{B_{t}}{\eta_{t}}$ such that $\eta_{t}\epsilon_{t}\leq B_{t}$, this leads to
$A_{t}$ $=\frac{\eta_{t}^{q_{\ast}}}{\alpha_{t}^{q_{\ast}-1}}=\Big{(}\frac{\eta_{t}}{\alpha_{t}}\Big{)}^{q_{\ast}}\alpha_{t}=\Big{(}\frac{t}{\tau(q-\beta)}\Big{)}^{\frac{q_{\star}}{q_{\star}-1}}\alpha_{t}$
$\propto\Big{(}\frac{t}{\tau}\Big{)}^{q}\cdot\frac{\alpha}{t^{\beta}}=\alpha t^{q-\beta}\cdot\frac{1}{\tau^{q}}$
$A_{t}$ $\propto\frac{\alpha}{\tau^{q}}t^{\frac{\kappa q+\kappa-q}{q}}$
@@
$F(x_{T+1}^{ag})-F_{\star}$ $=\mathcal{O}_{q,\kappa}\Big{(}\frac{\tau^{q}D_{\psi}(x_{\star},x_{1})}{\alpha T^{\frac{\kappa q+\kappa-q}{q}}}+\frac{\alpha^{\frac{\kappa}{q-\kappa}}L^{\frac{q}{q-\kappa}}}{\mu^{\frac{\kappa}{q-\kappa}}}\frac{\tau^{q}\log(T)}{T^{\frac{\kappa q+\kappa-q}{q}}}\Big{)},$
@@
Appendix D Proof for bounded iterate
Theorem D.1.
Let’s assume that $D_{\psi}(x_{\star},x_{1})\leq B$. Under assumption 1, we also suppose that we are running Algorithm 1 with parameters tuned such that there are base $K$ and exponents $n_{1},n_{2}\geq 0$ (possibly depending on $(L,\tau,\mu)$) ensuring $K^{n_{1}}>B$ and for all $t\geq 1$
$\max\left(\Big{(}\frac{qL\eta_{t}}{\kappa}\Big{)}^{\frac{1}{q-1}},\Big{(}\frac{L\alpha_{t}}{\kappa}\Big{)}^{\frac{1}{q-1}}\right)\leq K^{n_{1}}t^{n_{2}}.$
If we write $R_{t}=\max\{\|x_{t}-x_{\star}\|,\|x_{t}^{ag}-x_{\star}\|,\|x_{t}^{md}-x_{\star}\|\}$, then for all $t\geq 1$
$R_{t}=\mathcal{O}\Big{(}K^{\frac{n_{1}(q-1)}{q-\kappa}}t^{\frac{(q-1)(n_{2}+1)}{q-\kappa}}\Big{)}$
Proof.
We recall from our Algorithm 1 that
$x_{t+1}$ $=\arg\min_{x\in\mathbb{R}^{d}}\{\eta_{t}\langle\nabla F(x_{t}^{md}),x\rangle+D_{\psi}(x,x_{t})\}$
$x_{t+1}^{ag}$ $=\arg\min_{x\in\mathbb{R}^{d}}\{\alpha_{t}\langle\nabla F(x_{t}^{md}),x\rangle+\frac{\|x-x_{t}^{md}\|^{q}}{q}\}$
Since our setting is unconstrained, the (sub-)gradients have to cancel :
$\eta_{t}\nabla F(x_{t}^{md})+\nabla\psi(x_{t+1})-\nabla\psi(x_{t})=0$
$\alpha_{t}\nabla F(x_{t}^{md})+\phi_{q}(x_{t+1}^{ag}-x_{t}^{md})=0$
with $\psi(x)\propto\frac{\|x\|^{q}}{q}$ and $\phi_{q}(x)=\nabla\Big{(}\frac{\|x\|^{q}}{q}\Big{)}$. By the weak smoothness property of $F$:
$\|\nabla F(x_{t}^{md})\|_{\ast}\leq\frac{L}{\kappa}\|x_{t}^{md}-x_{\star}\|^{\kappa-1}.$
[…]
$\frac{\|x_{t+1}-x_{t}\|^{q}}{q}\leq D_{\psi}(x_{t},x_{t+1})+D_{\psi}(x_{t+1},x_{t})$ $=\langle\nabla\psi(x_{t+1})-\nabla\psi(x_{t}),x_{t+1}-x_{t}\rangle.$
Applying the optimality condition for $x_{t+1}$, this becomes
$\frac{\|x_{t+1}-x_{t}\|^{q}}{q}\leq\eta_{t}\langle\nabla F(x_{t}^{md}),x_{t}-x_{t+1}\rangle$
and by applying Cauchy-Schwartz, we obtain:
$\frac{\|x_{t+1}-x_{t}\|^{q}}{q}\leq\eta_{t}\|\nabla F(x_{t}^{md})\|_{\ast}\times\|x_{t}-x_{t+1}\|,$
[…]
$\|x_{t+1}^{ag}-x_{t}^{md}\|^{q}$ $=\langle\phi_{q}(x_{t+1}^{ag}-x_{t}^{md}),x_{t+1}^{ag}-x_{t}^{md}\rangle$
$\leq\alpha_{t}\|\nabla F(x_{t}^{md})\|_{\ast}\times\|x_{t+1}^{ag}-x_{t}^{md}\|.$
By plugging in these results in the previous conditions, we derive the following bounds:
$\|x_{t+1}-x_{t}\|^{q-1}\leq\frac{qL\eta_{t}}{\kappa}\|x_{t}^{md}-x_{\star}\|^{\kappa-1}$
$\|x_{t+1}^{ag}-x_{t}^{md}\|^{q-1}\leq\frac{L\alpha_{t}}{\kappa}\|x_{t}^{md}-x_{\star}\|^{\kappa-1}.$
From the triangle inequality, we then obtain :
$\|x_{t+1}-x_{\star}\|$ $\leq\|x_{t}-x_{\star}\|+\|x_{t+1}-x_{t}\|$
$\leq\|x_{t}-x_{\star}\|+\Big{(}\frac{qL\eta_{t}}{\kappa}\Big{)}^{\frac{1}{q-1}}\|x_{t}^{md}-x_{\star}\|^{\frac{\kappa-1}{q-1}}$
and
$\|x_{t+1}^{ag}-x_{\star}\|$ $\leq\|x_{t}^{md}-x_{\star}\|+\Big{(}\frac{L\alpha_{t}}{\kappa}\Big{)}^{\frac{1}{q-1}}\|x_{t}^{md}-x_{\star}\|^{\frac{\kappa-1}{q-1}}.$
Since $x_{t+1}^{md}$ is a middle point, we can also deduce that
$\|x_{t+1}^{md}-x_{\star}\|$ $=\|\lambda_{t+1}(x_{t+1}^{ag}-x_{\star})+(1-\lambda_{t+1})(x_{t+1}-x_{\star})\|$
$\leq\lambda_{t+1}\|x_{t+1}^{ag}-x_{\star}\|+(1-\lambda_{t+1})\|x_{t+1}-x_{\star}\|$
$\leq\max\Big{(}\|x_{t+1}^{ag}-x_{\star}\|,\|x_{t+1}-x_{\star}\|\Big{)}$
From the above inequality and using the assumption on $(\alpha_{t},\eta_{t})$ it follows that:
$R_{t+1}\leq R_{t}+K^{n_{1}}t^{n_{2}}R_{t}^{\frac{\kappa-1}{q-1}}.$
This recursive inequality shows that the growth is at most polynomial, implying that the sequence $R_{t}$ cannot grow faster than a polynomial rate.
We will show by induction that $R_{t}\leq K^{S_{t}}t^{M_{2}}$, where $M_{2}:=\max\Big{\{}\frac{(q-1)(n_{2}+1)}{q-\kappa},1\Big{\}}$ and $S_{t}:=n_{1}\sum_{i=0}^{t-1}\Big{(}\frac{\kappa-1}{q-1}\Big{)}^{i}$ is a geometric sum. First, we notice that
$n_{2}+1$ $\leq\frac{q-\kappa}{q-1}M_{2}$
$\implies n_{2}+M_{2}\frac{\kappa-1}{q-1}$ $\leq M_{2}-1$
$\implies t^{n_{2}+M_{2}\frac{\kappa-1}{q-1}}$ $\leq t^{M_{2}-1}\leq(t+1)^{M_{2}}-t^{M_{2}}.$
Then, by induction it then follows that
$R_{t+1}\leq R_{t}+K^{n_{1}}t^{n_{2}}R_{t}^{\frac{\kappa-1}{q-1}}$ $\leq K^{S_{t}}t^{M_{2}}+K^{n_{1}+S_{t}\times\frac{\kappa-1}{q-1}}t^{n_{2}+M_{2}\frac{\kappa-1}{q-1}}$
$\leq K^{S_{t+1}}\Big{(}t^{M_{2}}+t^{n_{1}+M_{2}\frac{\kappa-1}{q-1}}\Big{)}$
$\leq K^{S_{t+1}}(t+1)^{M_{2}}$
Since $S_{t}<\frac{n_{1}}{1-\frac{\kappa-1}{q-1}}=\frac{n_{1}(q-1)}{q-\kappa}$ , we can simplify $R_{t}=\mathcal{O}\Big{(}K^{\frac{n_{1}(q-1)}{q-\kappa}}t^{\frac{(q-1)(n_{2}+1)}{q-\kappa}}\Big{)}$. ∎
Appendix E Proof for binary search
Before proving Theorem 4.1, we make two quick observations.
Lemma E.1.